Journalist Report – December 29th

Journalist Report
Ben Durkee, Crew 236 Journalist

Sol 10

I’m starting to think that we’re the problem children of Mars. Under our watch, the Hab’s toilet, heater, and power system have independently acted up. I’m not superstitious, but I could be convinced that we’re cursed. Maybe I’m a little superstitious.

I won’t go into the details of the toilet issue – reliving it in my head triggers my gag reflex. Just know that these Martian boys are making do.

The heater has occasionally been blowing cold air whenever we look away from it, like a playful thermal Weeping Angel. Explains why everyone’s staterooms have been comfortable at night while the loft becomes a cryogenic chamber. I always thought that heat rises, but maybe physics work differently on Mars.

The sun has been playing a one-sided game of hide-and-seek, proliferating our power predicament primarily produced by a perturbing propane paucity. I am so sorry about that sentence. I think I blacked out for a few seconds – probably from thinking about the sewage again.

But at the same time, maybe this crew’s dysfunction is our strength! We powered down all non-necessary systems and had breakfast in the dark, brought together by the camaraderie of our new Amish lifestyle. It was kinda like when you were a kid and the power went out, so you gathered around a flashlight with your family and chowed down on the ice cream before it could melt. Man, I miss ice cream. Maybe my family too, just an ounce. But mostly ice cream.

We also had our first EVA composed of all Martian rookies today! And it went off without a hitch. Take that, curse!

In the meantime, we veterans had a day at home, in which the others helped me with my outreach. By outreach, I mean me belting out ABBA’s Mamma Mia in various locations across the station. It’s hard work, you know.

Eventually, the sun overcame its cowardice and we were able to return to business as usual. We played some card games in the Hab. Dylan showed us some gorgeous photos he’s taken of various celestial bodies with the observatory. We wound down with a delicious Filipino stew made by Chef Vlad and Chef Pavi.

Later, we’ll be celebrating our newfound electricity by finishing Apollo 13. We’ve been trying for a few nights in a row now, but always get rudely interrupted by that pesky curse. Third time’s the charm, I hope.

Oh, I know what the problem is! Mars must be in retrograde – that’s gotta be it. I’ll be speaking with the crew astronomer to get this remedied right away. I bet he’s a Taurus. Classic Taurus.

Operations Report – December 29th

Crew 236 Operations Report 29-DEC-2021

SOL: 10

Name of person filing report: Pavithra Ravi

Non-nominal systems: None


Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: Left unplugged after last EVA to save power. To be
charged by Shannon.

Opportunity rover used: Yes
Hours: 77.8
Beginning charge: 100%
Ending charge: 89%
Currently charging: Left unplugged after EVA to save power. To be
charged by Shannon.

Curiosity rover used: Yes
Hours: 157.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: Left unplugged after EVA to save power. To be
charged by Shannon.

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: To be charged by Shannon.

General notes and comments: None

Summary of Hab operations: The toilet flush is not functioning and
there is about an inch of waste stuck to the bottom of the tank that
we have not been able to flush away. In the past 24 hours, we have
flushed the toilet about 4 times (by filling buckets in the shower and
emptying it into the toilet), to no avail. Flushing the toilet for
individual use with a little bucket has not been too rough, so we can
make it work till the end of our rotation, after which a new toilet
will be installed.

We were in power conservation mode today due to cloud cover limiting
the available solar power.


Water (static tank): approx. 100-150 gallons after all that flushing!
Water (loft tank): 21 gallons
Water Meter: 01542759 units

Static to Loft Pump used – yes or no: Yes. It was filled from 32 to 55
gallons this afternoon.
Static tank pipe heater (on or off): On
Static tank heater (on or off) Off

Toilet tank emptied (no or yes): yes, around 14:30

Summary of internet: Nominal

Summary of suits and radios: Nominal

We ran out of one can of plexiglass cleaner and are onto a second one.

Summary of GreenHab operations: Nominal

WATER USE: 7 gallons were used today. This leaves about 158 gallons
in the tank.

Heater (on or off): On

Supplemental light (on or off): On

Harvest: 5g of parsley, 2g of wheatgrass, 1g Thai hot pepper

Summary of ScienceDome operations: Nominal

Dual split, Heat or AC (on or off): off

Summary of RAM operations: Tools used, work done

Nothing used.

Summary of any observatory issues: Nothing new to report.

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: None

Sol Summary – December 29th

Crew 236 Sol Summary Report 29-DEC-2021

Sol: 10

Summary Title: Martian Twilight Zone

Author’s name: Kasey Hilton

Mission Status: Nominal, last day of Sim tomorrow!

Sol Activity Summary: After a late night of battling with emptying the toilet tank, the crew had a late wake-up and a delicious pancake breakfast. We quickly learned that we needed to go into energy conservation mode since the solar panels couldn’t charge while it was cloudy and the propane level for the generator was low.

That meant we spent the afternoon working on research and outreach until our laptop batteries died and playing more Durak in the window lit Hab. Today’s EVA was scheduled for the second half of the afternoon, so the remaining crew continued work on research, outreach, and troubleshooting the toilet issues. And by “outreach”, I mean our Crew Journalist, Ben, filming himself belting out Mama Mia in the Hab living room.

During a quick evening conversation with The Outpost, we had a good laugh at the idea that the last few days have felt like a Martian version of The Twilight Zone with a handful of things breaking or not working how they should (read the Operations report, you’ll understand). After trying out a new Philipino stew for dinner, we will attempt to end the night with Apollo 13… again.

Look Ahead Plan: Tomorrow is our last day of Sim, so some of the crew members will be going on the last EVA of the rotation and the remaining crew will be wrapping up the last parts of their research and outreach projects in the Hab.

Anomalies in work: None

Weather: Partly cloudy, temperatures in the high 30s/low 40s F

Crew Physical Status: Nominal (if not a bit hyperactive)

EVA: Today’s EVA to Barianca Butte was successful. Vladimir (GHO) scouted and found three additional Hab locations, while Tyler (SCI) generated a second heat map of the Barianca Butte with different weather conditions

Reports to be filed: Operations Report, GreenHab Report, EVA Report, EVA Request, Health and Safety Report, Sol Summary, and Journalist Report (with photos)

Support Requested: None

Research Report – December 29th

Astronomy Report

Name: Dylan Dilger
Crew: 236
Date: 12/29/2021


Robotic Telescope Requested (choose one): Montana Learning Center

Objects Viewed: IC434 (Horsehead Nebula), M95, M96, NGC2244 (Rosette Nebula), M51, M42, M78, Almach
Problems Encountered: None

Solar Features Observed: None
Problems Encountered: None

Greenhab Report – December 29th

GreenHab Officer: Vladimir Zeltsman

Environmental Control: Around 17% Humidity
Average Temperatures: 78degrees
Hours of Supplemental Light: 7 hours
Daily Water Usage For Crops: 7 gallons
Daily Water Usage For Research and/or Other Purposes: None
Water In Blue Tank: 158 Gallons
Times Of Watering For Crops: 9:30 am
Changes to Crops: None
Narrative: None
Harvest: 2g Wheat Grass, 1g Thai hot peppers, 5g parsley
Support/supplies needed: None

Copyright © The Mars Society. All rights reserved. | Main Site